DEAR SIR , 1 . Two objects each of mass 1 . 5 kg , are moving in the same straight time but in opposite dirctions with a velocity of 2.5 m /s . they collide with each other and stick to each other . calculate the velocity of combined object after the collision .

DEAR  SIR ,  
 
     1  .  Two objects each of mass 1 . 5 kg , are moving in the same straight time but in opposite  dirctions with a velocity of 2.5 m /s . they collide with each other and stick to each other . calculate the velocity of combined object after the collision . 

Grade:9

1 Answers

Nisarga Rajendra Khairnar
9 years ago
The velocity of combined object will be 0 m/s. Because the momentum of both the objects is same but in opposite direction therefore it will cancel each other 
momentum of 1st particle = m X v = 1.5 X 2.5 = 3.75 kgm/s
momentum of 2nd particle =m X v = 1.5 X -2.5=-3.75kgm/s
therefore the resultant momentum = 3.75 +(-3.75) = 0 kgm/s
as momentum = m X v mass of the combined particle = 2X1.5=3kg
therefore 0kgm/s=3 X v   hence velovity=0m/s